OpenStudy (anonymous):

7(x − 1) = 5(x + 2) + 1

OpenStudy (uri):

Solve the bracket first 7(x − 1) = 5(x + 2) + 1 7x-7=5x+10+1 Now add 7 both sides 7x-7=5x+10+1 +7 +7 ========= 7x=5x+17+1

OpenStudy (uri):

7x=5x+17+1 move the 5x on the left side so 7x-5x=17+1 2x=18

OpenStudy (uri):

Divide 18 with 2 x=18/2 x=9

OpenStudy (anonymous):

Thanks

OpenStudy (uri):

:)